OpenStudy (lena772):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = one-halfx + (x - 6) What is the regular price of a child's ticket? $50 $56 $75 $81

OpenStudy (lena772):

\[78=\frac{ 1 }{ 2 }x+(x-6)\]

OpenStudy (lena772):

@ganeshie8 MY FAVORITE PERSON! :)

ganeshie8 (ganeshie8):

hey :)

OpenStudy (lena772):

heyy :)

ganeshie8 (ganeshie8):

\(78=\frac{ 1 }{ 2 }x+(x-6) \) multiply left and right wid 2 \(78*2 =2*\frac{ 1 }{ 2 }x+2*(x-6) \) \(78*2 =x+2x-12 \) simplify and solve x first

OpenStudy (lena772):

156=3x-12 156+12=3x 168=3x 59=x

OpenStudy (lena772):

I mean 56=x

ganeshie8 (ganeshie8):

Yes ! but \(x\) is \(adult\) ticket here in the given equaiton

ganeshie8 (ganeshie8):

child ticket = adult - 6

OpenStudy (lena772):

I'm sorry :(

ganeshie8 (ganeshie8):

adult ticket = 56 child ticket = 56-6 = 50

OpenStudy (lena772):

Ohh! Ok my bad. Cuz it says she paid 78 for an adult ticket and two children's tickets

ganeshie8 (ganeshie8):

OMG ! you're right !! i missed that completely lol wait lets work it again

ganeshie8 (ganeshie8):

hey no, actually 50 is correct, the question is asking for \(regular\) price of child ticket oly

OpenStudy (lena772):

No you're right. I didn't realize she bought them half price.

ganeshie8 (ganeshie8):

Yes 78 is after discount... you gave me heart attack for a sec :o

OpenStudy (lena772):

Lol sorry my fault!

ganeshie8 (ganeshie8):

np :))